2. A man who has a mass of 80 kg can run up a flight of stairs of 20 steps in 5 seconds. Given the each steps in 20cm high, how much power does the man use to run up the flights of stairs?


1
Expert's answer
2022-03-02T14:38:51-0500

By definition, the power is:


"P = \dfrac{W}{t}"

where "W" is the work done by the man, and "t = 5s" is the time. In order to run up the stairs the man should do the following work:


"W = mgh"

where "m = 80kg" is the mass of the man, "g = 9.8m\/s^2" is the gravitational acceleration, and "h = 20\cdot 20cm = 4m" is the total height of the stairs. Thus, obtain;


"P = \dfrac{mgh}{t} = \dfrac{80\cdot 9.8\cdot 4m}{5s} \approx 627W"

Answer. 627W